OpenStudy (anonymous):

i need to factor 6x(x-1)=63-19x

OpenStudy (anonymous):

simplify first, \[6x^{2}-6x =63-19x\] then gather all values on one side of the equation \[6x ^{2}-6x +19x -63=0\] combine like terms \[6x ^{2}+13x -63=0\] then factor, and you end up with \[(3x-7)(2x+9)=0\]

OpenStudy (anonymous):

thanks alot.. i got up until 6x^2+13x-63=0 but i couldn't find any factors.